Five identical point objects of equal masses M each are located at the five vertices of a regular hexagon of side a . The magnitude of the gravitational field at the centre of the hexagon is 
 
A. 0
 
B. 5GM/a^2
 
C.GM/a^2
 
D.6GM/a^2
 

Grade:11

1 Answers

Atul Nagar
5 years ago
Let us consider the Hexagon with vertices A,B,C,D,E and F and mass M is stationed at vertices A to A to E. The mass as A and D creates the field which are equal and opposite and hence cancel each other. Similarily the mass at B and E creates field which are equal in magnitude but opposite in direction and hence cancels out.
Thus total field at Center of Hexagon = Ea+Eb+Ec+Ed+Ef
Net feild at Centre = GM/a^2+GM/a^2+GM/a^2-GM/a^2-GM/a^2 = GM/A^2 i.e option C.